OpenStudy (anonymous):

Let c be any constant number. Which of the following will always be perpendicular to −3x + y = 2?

OpenStudy (amistre64):

depends on the following i spose; but perps swap coeffs and negate one of them

OpenStudy (amistre64):

−3x + y = 2 −3x + y = c x -3y = 2 x + 3y = 2 whould do it

OpenStudy (amistre64):

i forgot to edit the 2 into a c all the way thru, but you get the point

OpenStudy (anonymous):

thank yu